OpenStudy (anonymous):

Geogebra said this function is: 8.63 e^(-0.53x) F(x)=8.63e^(-0.53x) Y=8.63e^(-0.53x) In y/8.63= In e^(-0.53x) In y – In 8.63= In e^(-0.53x) In y= -0.53x + 2.16 Then it said change it into half-life function but I don't get how I am supposed to write a half life function out of that Half life formula: A=[1/2]^x

OpenStudy (phi):

logs are confusing but you could say this: e to some power = 1/2 by definition, this is \[ e^{ln(1/2)} = \frac{1}{2} \] if you raise both sides to the 1/ln(1/2) \[ e^{\frac{ln(1/2)}{\ln(1/2)}} = \left(\frac{1}{2}\right)^{\frac{1}{\ln(1/2)}} \] this last step used the idea that \[ (e^x)^\frac{1}{x} = e^{\frac{x}{x}} = e^1= e\] so you now know \[ \left(\frac{1}{2}\right)^{\frac{1}{ln(1/2)}} = e\] plot that mess into your equation instead of e: \[ F(x)=8.63e^{-0.53x} \\ F(x)=\left(\left(\frac{1}{2}\right)^{\frac{1}{\ln(1/2)}}\right)^{-0.53x} \\ F(x)=\left(\frac{1}{2}\right)^{- \frac{0.53}{\ln(1/2)}x} \]

OpenStudy (phi):

**8.63 should still be in the last 2 equations

OpenStudy (anonymous):

thankss:) @phi